OpenStudy (anonymous):

Please help!!!!!! How do you solve this??? limx→4=(√x)-2/x−4

OpenStudy (hunus):

The function\[\sqrt{x}-\frac{2}{x}−4\] is continuous at x=4 so you can just plug in 4 everywhere you see x

OpenStudy (anonymous):

Sorry -- the actual equation looks like this\[\lim_{x \rightarrow 4} = \frac{ \sqrt{x}-2 }{ x-4 }\]

OpenStudy (kenljw):

The result is 0/0 therefore use L'Hospital Rule N' = 1/2x D' = 1 N'/D' = 1/2x as x goes to 4, 1/8

OpenStudy (anonymous):

thanks!